OpenStudy (hanihani):

How do I find the dimensions of a rectangle where the length is is bigger than the width by 1 centimeter. The area is 360 centimeters

OpenStudy (steve816):

First, note that length times width = Area

OpenStudy (hanihani):

yeah i get that

OpenStudy (hanihani):

I have to use a formula similar to (x+1)(x=1)=y

OpenStudy (steve816):

So then, solve this equation\[\Large x(x+1) = 360\]

OpenStudy (hanihani):

idk

OpenStudy (steve816):

Expand the equation to get it into ax^2+bx+c = 0 form. That would be\[\Large x^2+x-360=0\]

OpenStudy (steve816):

Now, use the quadratic formula to solve for x